If 16.0 g of Mg (OH)2 and 11.0 g of HCI are combined...what is the limiting reactant and how many grams of Mg CI2 will be produced?
1
Expert's answer
2020-11-23T06:59:03-0500

The equation is

Mg(OH)2 + 2HCl --> MgCl2 + 2H2O


"n_{Mg(OH)_2}=\frac{16.0g}{58.32g/mol}=0.274 mol"


"n_{HCl}=\frac{11.0g}{36.46g/mol}=0.302mol"


According to the equation, the mole ratio HCl : Mg(OH)2 = 2 : 1

"\frac{0.302}{0.274}<\frac{2}{1}" , hence HCl is the limiting reactant.


"m_{MgCl_2}=0.302mol_{HCl}\times\frac{1 mol_{MgCl_2}}{2mol_{HCl}}\times\frac{95.21g_{MgCl_2}}{1mol_{MgCl_2}}=14.4g"


Answer: HCl; 14.4 g